OpenStudy (anonymous):

Given the initial condition y(1)=4, find the particular solution of the equation y'=y(2x+3)

OpenStudy (anonymous):

I've solved it up to here: Dy/dx=y(2x+3) F(x)=2x+3 G(y)=y Integral(F(x))=Integral(1/G(y)) X^2+3x=Ln(|y|)+C_1

OpenStudy (loser66):

backward lny = x^2+3x +C

OpenStudy (anonymous):

Oh, the C is on the left side?

OpenStudy (anonymous):

Ok, so

OpenStudy (anonymous):

Ln(|y|)+C=X^2+3x

OpenStudy (anonymous):

Now we solve for C?

OpenStudy (anonymous):

Oh, I think that's been my mistake. I haven't been solving C last.

OpenStudy (anonymous):

y=e^(x^2+3x-C) 2=e^(1^2+3*1-C)

OpenStudy (loser66):

you can solve for C there, since you don't have particular solution. :)

OpenStudy (anonymous):

oh, found another mistake. y should equal 4 in the intial solution

OpenStudy (anonymous):

It seems that when I solve for C, I get 4-2*ln(2). Equation I solved: 4=e^(1^2+3*1-C)

OpenStudy (loser66):

\(lny = x^2+3x+C\\y = e^{x^2+3x+C}= C e^{x^2+3x}\\C = \dfrac{y}{e^{x^2+3x}}\)

OpenStudy (anonymous):

I have the initial condition though

OpenStudy (loser66):

now, plug x =1,y =4 in to get C

OpenStudy (anonymous):

~0.699?

OpenStudy (loser66):

I don't know!! to me, I let it as it is 4e^(-4)

OpenStudy (anonymous):

My choices for the particular solution are y=e^(x^2+3x) - 0.073 y=0.073e^(x^2+3x) y=e^(x^2+3x+0.073) y=e^(x^2+3x)+0.073

OpenStudy (anonymous):

Oh, I must've messed up when solving for C. How did you get 4/e^4?

OpenStudy (loser66):